# Too Many Frames, Not All Useful: Efficient Strategies for Long-Form Video QA

### v3

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, IntentQA, NExT-QA. The code can be found at https://github.com/jongwoopark7978/LVNet
@@ 1 Introduction
Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], IntentQA[22], and NExT-QA[41] as described in 4.2. This demonstrates the strong performance and general applicability of our approach.
@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], IntentQA [22], and NExT-QA [41]. The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. The second dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.). The third dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos.
@@
Table 1: Combined Results on IntentQA [22], EgoSchema [24], and NExT-QA [41]. We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits. The models are ordered based on the number of captions processed before answering the question. LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions. Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
Model IntentQA EgoSchema NExT-QA
Cap. Why? (%) How? (%) B./A. (%) All (%) Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%)
Vamos [36] - - - - - - 48.3 - - - - -
IG-VLM [17] - - - - 65.3 - 59.8 - 69.8 63.6 74.7 68.6
VideoLLaMA 2 [5] - - - - - - 53.3 - - - - -
InternVideo2 [38] - - - - - - 60.2 - - - - -
Tarsier [35] - - - - - - 61.7 - - - - 79.2
VIOLET [10] - - - - - 5 19.9 - - - - -
mPLUG-Owl [46] - - - - - 5 31.1 - - - - -
VideoAgent [37] - - - - - 8.4 54.1 8.2 72.7 64.5 81.1 71.3
MVU [30] - - - - - 16 37.6 16 55.4 48.1 64.1 55.2
MoReVQA [25] - - - - - 30 51.7 16 70.2 64.6 - 69.2
TraveLER [37] - - - - - - - $\sim$25 70.0 60.5 78.2 68.2
VFC [26] - - - - - - - 32 45.4 51.6 64.1 51.5
SeViLA [48] 32 - - - 60.9 32 22.7 32 61.3 61.5 75.6 63.6
ProViQ [6] - - - - - 60 57.1 60 - - - 64.6
VideoTree [40] $\leq$56 - - - 66.9 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5
FrozenBiLM [45] - - - - - 90 26.9 - - - - -
LifelongMemory [39] - - - - - 90 62.1 - - - - -
LangRepo [16] 90 62.8 62.4 47.8 59.1 180 41.2 90 64.4 51.4 69.1 60.9
LLoVi [52] 90 68.4 67.4 51.1 64.0 180 50.3 90 69.5 61.0 75.6 67.7
LVNet (ours) 12 75.0 74.4 62.1 71.7 12 61.1 12 75.0 65.5 81.5 72.9
@@ 4.2 Evaluation
We evaluate LVNet on the IntentQA, EgoSchema, and NExT-QA dataset and present our results in Table 1. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%. To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset. The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
We next evaluate on the IntentQA[22] dataset. Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
We finally evaluate on the NExT-QA[41] dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.

### v4

@@ Abstract
Long-form videos that span across wide temporal intervals are highly information redundant and contain multiple distinct events or entities that are often loosely related. Therefore, when performing long-form video question answering (LVQA), all information necessary to generate a correct response can often be contained within a small subset of frames. Recent literature explore the use of large language models (LLMs) in LVQA benchmarks, achieving exceptional performance, while relying on vision language models (VLMs) to convert all visual content within videos into natural language. Such VLMs often independently caption a large number of frames uniformly sampled from long videos, which is not efficient and can mostly be redundant. Questioning these decision choices, we explore optimal strategies for key-frame selection that can significantly reduce these redundancies, namely Hierarchical Keyframe Selector. Our proposed framework, LVNet, achieves state-of-the-art performance at a comparable caption scale across three benchmark LVQA datasets: EgoSchema, NExT-QA, IntentQA. The code can be found at https://github.com/jongwoopark7978/LVNet
@@ 1 Introduction
Our proposed LVNet achieves state-of-the-art results compared to the models that utilize a similar number of captions on three long-form video question answering benchmarks—EgoSchema[24], NExT-QA[41], and IntentQA[22] as described in 4.2. This demonstrates the strong performance and general applicability of our approach.
@@ 4.1 Experimental Setup
Datasets: Given the training free nature of our framework, we do not utilize any video datasets for training. Datasets are used purely for evaluation. We select three benchmark video visual question answering datasets focused on long-form videos for this purpose: EgoSchema [24], NExT-QA [41], and IntentQA [22]. The first dataset, EgoSchema, is a very long-form video question-answering dataset with 5031 questions. For each question, EgoSchema requires the correct answer to be selected between five given options based on a three-minute-long video clip. The second dataset, NExT-QA, is another rigorously designed video question answering benchmark containing questions that require causal & temporal action reasoning, and common scene comprehension to correctly answer. These questions are further classified as Causal (Cau.), Temporal (Tem.), and Descriptive (Des.) and we evaluate on its validation set containing 4996 questions over 570 videos. The third dataset, IntentQA [22] is based on NExT-QA videos corresponding to temporal and causal reasoning quetions. It consists of 16k multiple-choice questions which are classified as Why?, How? or Before/After (B./A.).
@@
Table 1: Combined Results on EgoSchema [24], NExT-QA [41], and IntentQA [22]. We compare LVNet against prior zero-shot models across three datasets, highlighting different task splits. The models are ordered based on the number of captions processed before answering the question. LVNet achieves state-of-the-art accuracies of 71.7%, 61.1%, and 72.9% on the three datasets, respectively, using just 12 frames compared to the models using the similar number of captions. Models with video-caption pretraining or utilizing significantly more captions than the 12 frames used by LVNet are de-emphasized in grey or downplayed in light green to ensure fairness with image-level pretraining or highlight caption efficiency.
Model EgoSchema NExT-QA IntentQA
Cap. Acc. (%) Cap. Cau. (%) Tem. (%) Des. (%) All (%) Cap. Why? (%) How? (%) B./A. (%) All (%)
Vamos [36] - 48.3 - - - - - - - - - -
IG-VLM [17] - 59.8 - 69.8 63.6 74.7 68.6 - - - - 65.3
VideoLLaMA 2 [5] - 53.3 - - - - - - - - - -
InternVideo2 [38] - 60.2 - - - - - - - - - -
Tarsier [35] - 61.7 - - - - 79.2 - - - - -
VIOLET [10] 5 19.9 - - - - - - - - - -
mPLUG-Owl [46] 5 31.1 - - - - - - - - - -
VideoAgent [37] 8.4 54.1 8.2 72.7 64.5 81.1 71.3 - - - - -
MVU [30] 16 37.6 16 55.4 48.1 64.1 55.2 - - - - -
MoReVQA [25] 30 51.7 16 70.2 64.6 - 69.2 - - - - -
TraveLER [37] - - $\sim$25 70.0 60.5 78.2 68.2 - - - - -
VFC [26] - - 32 45.4 51.6 64.1 51.5 - - - - -
SeViLA [48] 32 22.7 32 61.3 61.5 75.6 63.6 32 - - - 60.9
ProViQ [6] 60 57.1 60 - - - 64.6 - - - - -
VideoTree [40] 63.2 61.1 $\leq$56 75.2 67.0 81.3 73.5 $\leq$56 - - - 66.9
FrozenBiLM [45] 90 26.9 - - - - - - - - - -
LifelongMemory [39] 90 62.1 - - - - - - - - - -
LangRepo [16] 180 41.2 90 64.4 51.4 69.1 60.9 90 62.8 62.4 47.8 59.1
LLoVi [52] 180 50.3 90 69.5 61.0 75.6 67.7 90 68.4 67.4 51.1 64.0
LVNet (ours) 12 61.1 12 75.0 65.5 81.5 72.9 12 75.0 74.4 62.1 71.7
@@ 4.2 Evaluation
We evaluate LVNet on the EgoSchema, NExT-QA, and IntentQA dataset and present our results in Table 1. Models with video-caption pretraining are de-emphasized in grey to ensure fairness with image-level pertaining. Models utilizing significantly more captions than the 12 frames are downplayed in light green to consider caption efficiency. For EgoSchema, we achieve 61.1% on the fullest, the highest among the models utilizing approximately 12 captions. This result outperforms VideoAgent, the next best model using 8.4 captions, by 7%. To ensure a fair comparison, we ablate the number of captions for LVNet in LABEL:tab:ablation:numberofframecaptions, observing that LVNet with 8 captions still outperforms VideoAgent with 8.4 captions by 4.2% in the EgoSchema subset. The superior accuracy of LVNet over other keyframe selectors, such as those used in VideoAgent and TraveLER, is discussed in Section 3.4.
We next evaluate on the NExT-QA[41] dataset. This dataset has a particular focus on both temporal and casual reasoning based question-answer pairs. Our approach achives state-of-the-art performance on this benchmark outperforming prior work among the models utilizing approximately 12 captions. In fact, our LVNet with just 12 frame captions achieves 72.9% overall accuracy, outperforming VideoAgent [37] (71.3% at 8.2 captions) by 1.6%.
We finally evaluate on the IntentQA[22] dataset. Our approach achieves state-of-the-art results on this benchmark, outperforming all prior work, including the de-emphasized models with video-caption pretraining and the downplayed models utilizing significantly more captions than 12 frames. In fact, LVNet shows a substantial improvement (+4.8%) over the next best model, VideoTree [40] (66.9% with 90 captions), achieving 71.7% accuracy with only 12 frame captions.
@@ 5 Conclusion
Acknowledgements:
This work was supported by Electronics and Telecommunications Research Institute (ETRI) grant funded by the Korean government [24ZB1200, Research of Human-centered autonomous intelligence system original technology]. This work was also supported by the Institute of Information & communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2024-00336738, Development of Complex Task Planning Technologies for Autonomous Agents, 100%)